OpenStudy (anonymous):

Simplify 64 to the negative 1 third power. −4 negative 1 over 4 1 over 4 4

OpenStudy (freckles):

\[64^\frac{-1}{3}=\frac{1}{64^\frac{1}{3}}\]

OpenStudy (freckles):

what number cube equals 64?

OpenStudy (freckles):



OpenStudy (anonymous):

sorry the wifi went off and can you explain it again

OpenStudy (anonymous):

oh lol well 4^3 = 64

OpenStudy (freckles):

so since 4^3=64 then 64^(1/3)=4

OpenStudy (freckles):